# Diffusion2: Dynamic 3D Content Generation via Score Composition of Orthogonal Diffusion Models

### v2

@@ Abstract
Recent advancements in 3D generation are predominantly propelled by improvements in 3D-aware image diffusion models which are pretrained on Internet-scale image data and fine-tuned on massive 3D data, offering the capability of producing highly consistent multi-view images. However, due to the scarcity of synchronized multi-view video data, it is impractical to adapt this paradigm to 4D generation directly. Despite that, the available video and 3D data are adequate for training video and multi-view diffusion models that can provide satisfactory dynamic and geometric priors respectively. In this paper, we present Diffusion2, a novel framework for dynamic 3D content creation that leverages the knowledge about geometric consistency and temporal smoothness from these models to directly sample dense multi-view and multi-frame images which can be employed to optimize continuous 4D representation. Specifically, we design a simple yet effective denoising strategy via score composition of video and multi-view diffusion models based on the probability structure of the images to be generated. Owing to the high parallelism of the image generation and the efficiency of the modern 4D reconstruction pipeline, our framework can generate 4D content within few minutes. Furthermore, our method circumvents the reliance on 4D data, thereby having the potential to benefit from the scalability of the foundation video and multi-view diffusion models. Extensive experiments demonstrate the efficacy of our proposed framework and its capability to flexibly adapt to various types of prompts.
@@ 1 Introduction
Figure 1: Diffusion2 is designed to generate dynamic content by creating a dense multi-frame multi-view image matrix in a highly parallel denoising diffusion process with the combination of the foundation video diffusion model and multi-view diffusion model. The generated image can be used to construct a full 4D representation by being fed into an off-the-shelf 4D reconstruction pipeline.
@@
Spurred by the advances from generative image models ho2020denoising ; song2020denoising ; song2020score ; karras2022elucidating ; zhang2023adding , automatic 3D content creation poole2022dreamfusion ; wang2024prolificdreamer ; tang2023dreamgaussian ; hong2023lrm has witnessed remarkable progress in terms of efficiency, fidelity, diversity, and controllability. Coupled with the breakthroughs in 4D representation yang2023deformable ; wu20234d , these advancements further foster substantial development in dynamic content (4D) generation singer2023text ; bahmani20234d ; jiang2023consistent4d ; zhao2023animate124 ; ren2023dreamgaussian4d ; gao2024gaussianflow , which holds significant value across a wide range of applications in animation, film, game, and MetaVerse.
@@
Recently, 3D content generation has achieved considerable breakthroughs in efficiency. Some works liu2023zero ; liu2023syncdreamer ; shi2023mvdream ; wang2023imagedream ; tang2024mvdiffusion++ inject stereo knowledge into the image generation model, enabling these 3D-aware image generators to produce consistent multi-view images, thereby effectively stabilizing and accelerating the optimization. Other efforts hong2023lrm ; chen2024v3d ; voleti2024sv3d ; zuo2024videomv ; tang2024lgm attempt to directly generate 3D representations, such as triplane chen2022tensorf or Gaussian splatting kerbl3Dgaussians . However, the efficiency improvement from these works is largely data-driven wu2023omniobject3d ; reizenstein2021common ; yu2023mvimgnet ; deitke2023objaverse . Consequently, it is infeasible to adapt these approaches to 4D generation due to the scarcity of synchronized multi-view video data. Therefore, most existing optimization-based 4D generation works still suffer from slow and unstable optimization.
@@
However, despite the paucity of 4D data, there are vast available monocular video data and static multi-view image data. Existing works have demonstrated that it is feasible to train diffusion-based generative models to learn the distribution of these two classes of data separately voleti2024sv3d ; liu2023syncdreamer ; tang2024mvdiffusion++ ; blattmann2023align ; blattmann2023stable . Considering that video diffusion model stores the prior of motion and temporal smoothness, and multi-view diffusion model has sound knowledge of geometrical consistency, combining the two generative models to generate 4D assets becomes a highly promising and appealing approach.
@@
Therefore, in this paper we propose a novel 4D generation framework, which combines both the video diffusion model and multi-view diffusion model to directly sample multi-frame multi-view image array imitating the photographing process of 4D content. To demonstrate how we realize this combination, we assume that such an image matrix has a nice structure: the elements not in the same row and column are conditionally independent of each other. Based on the property, we design a simple yet effective denoising strategy in which the estimated score is just the convex combination of the scores predicted by two foundation diffusion models. Our formulation is easy to be adapted to various prompts including single image, single-view video, and static 3D content. Unlike the existing optimization-based counterparts, our image generation is highly parallel. Combined with efficient modern 4D reconstruction methods, we can generate high-fidelity and diverse 4D assets within just several minutes. Besides, our approach can also potentially benefit from the further development of foundation diffusion models videoworldsimulators2024 .
@@
Our contributions can be summarized into threefold: (i) We present a novel 4D content generation framework that achieves zero-shot generation of realistic multi-view multi-frame image arrays, which can be integrated with an off-the-shelf modern 4D reconstruct pipeline to efficiently create 4D content. (ii) We identify the conditional independence that existed in the distribution of the elements composing the image arrays. And based on it, we design a simple yet effective joint denoising strategy that combines video diffusion model and multi-view diffusion model to directly sample multi-view multi-frame image arrays from their natural distribution. (iii) Systematic experiments demonstrate that our proposed method can achieve satisfactory results under different types of prompts, including single image, single-view video, and static 3D content.
2 Reletaed work
3D generation
@@ 2 Reletaed work
3D generation aims at creating static 3D content from different prompts like text or image. Early efforts employed GAN-based approaches gao2022get3d ; schwarz2020graf . Recently, significant breakthroughs have been achieved alongside the emergence of diffusion models ho2020denoising in this domain. DreamFusion poole2022dreamfusion introduced score distillation sampling (SDS) to unleash the creativity in diffusion models. Although such approach has exhibited promising results, the original form of SDS encounters challenges such as mode collapse, multi-face Janus issues, and the slow optimization. A series of subsequent works wang2024prolificdreamer ; shi2023mvdream ; wang2023imagedream ; pan2024enhancing ; tang2023dreamgaussian ; yi2023gaussiandreamer try to address these problems by modifying this mechanism. On the other hand, some studies nichol2022point ; jun2023shap ; hong2023lrm ; tang2024lgm ; wang2024crm have explored the direct generation of 3D representations using diffusion models. Another line of research liu2023zero ; liu2023syncdreamer ; long2023wonder3d ; chen2024v3d ; tang2024mvdiffusion++ focuses on generating dense multi-view images with sufficient 3D consistency by training or fine-tuning 2D diffusion models on 3D datasets to make them more suitable for 3D generation tasks. The generated images can be used for reconstruction to obtain textured meshes, point clouds or implicit radiance fields. We also adopt this approach of directly generating consistent images for reconstruction. But unlike these 3D counterparts, there is no large-scale multi-view synchronized video data. Therefore, we opt to combine geometrical consistency priors and video dynamic priors to generate images.
Video generation
@@
Video generation and prediction is an active field that has gained increasing popularity. Recent diffusion model-based video generation methods have exhibited unprecedented levels of realism, diversity, and controllability. Notably, recent breakthroughs have demonstrated the scalability of video generation models combined with Transformers and their potential as physical world simulators. Video LDM blattmann2023align was among the first works to apply the latent diffusion framework rombach2022high to video generation. The subsequent work SVD blattmann2023stable followed its architecture and made effective improvements to the training recipe. W.A.L.T gupta2023photorealistic employed a transformer with window attention tailored for spatiotemporal generative modeling to generate high-resolution videos. VDT lu2023vdt introduced the video diffusion transformer to flexibly capture long-distance spatiotemporal context in videos and a spatial-temporal mask mechanism to uniformly handle different video generation tasks. The recently introduced SORA videoworldsimulators2024 demonstrated a remarkable capability to generate arbitrarily sized long videos with intuitively physical fidelity. Models trained on large-scale video data can generate videos with consistent geometry and realistic dynamics. Furthermore, video diffusion models can also be considered as effective 3D generators chen2024v3d ; han2024vfusion3d ; voleti2024sv3d to generate multi-view consistent images. Therefore, we build our method on this flourishing domain.
4D generation
@@
Animating category-agnostic stuff is a challenging problem and has been receiving a lot of attention from both academia and industry. Compared to 3D generation, 4D generation requires not only predicting consistent geometry but also generating realistic and diverse dynamics. Recent works on 4D generation could be classified into two main streams based on the type of input prompt. The first class of methods predicts 4D representations given a single image and text description as input. For instance, MAV3D singer2023text directly deploys SDS into 4D generation and proposes a three-stage training scheme to stably generate high-resolution videos. DreamGaussian4D ren2023dreamgaussian4d adopt a similar three-stage training scheme but switches the underlying 4D representation from Hexplane cao2023hexplane to the deformable 3D Gaussian and replaces the third stage with the refinement of high-resolution UV texture maps, thereby achieving significantly efficiency improvement. 4DGen yin20234dgen first generates a set of candidate 3D assets according to the image or text prompt and then grounds the 4D generation on user-specified static 3D assets and monocular video sequences. 4D-fy bahmani20234d also adopts the idea of combining diffusion priors. But different from our method, it combines them during SDS, resulting in an extremely slow generation process that typically takes several days. Another line of work predicts dynamic objects from a single-view videos. Although the motion is largely dictated under this setting, generating spatiotemporal consistent structure still entails considerable uncertainty due to the limited view. Therefore, Consistent4D jiang2023consistent4d is first proposes to use a generation approach to address this task. Subsequently, Efficient4D pan2024fast mimics a photogrammetry-based 3D object capture pipeline by directly generating multi-frame multi-view captures of dynamic 3D objects through SyncDreamer-T and reconstructing 4D representations with them. Note that although Efficient4D shares the same SDS-free philosophy as ours, it possesses a strong architecture bias of the foundation diffusion model and is incapable of synthesizing novel dynamics. Compared to previous works, our framework can efficiently generate diverse dynamic 4D content, avoiding the slow, unstable, and intricate multi-stage optimization, and has the potential to continuously benefit from the scalability of the underlying diffusion model.
@@ 3 Method
Figure 2: The overall pipeline of Diffusion2. (i) Given a reference image, Diffusion2 first independently generates the animation under the reference view (denoted $\mathcal{I}_{1:V,0}$) and the multi-view images at the reference time (denoted $\mathcal{I}_{0,1:F}$) as the condition for the subsequent generation of the full matrix, denoted $\mathcal{I}$. Depending on the type of given prompt, the condition images $\mathcal{I}_{1:V,0}$ or $\mathcal{I}_{0,1:F}$ can be specified by users. (ii) Then, Diffusion2 directly samples a dense multi-frame multi-view image array by blending the estimated scores with a weighting scheduler from pretrained video and multi-view diffusion models in the reverse-time SDE. (iii) The generated image arrays are employed as supervision to optimize a continuous 4D content representation.
In this section, we present a novel framework designed for efficient and scalable generation of 4D content, which consists of two stages as depicted in Figure 2. In Section 3.1 (stage-1), we will discuss how to generate a dense multi-view multi-frame image array for reconstruction through a highly parallelizable denoising process by integrating the pretrained video diffusion model and multi-view diffusion model, and why it is feasible. In Section 3.2 (stage-2), we will briefly illustrate how to robustly reconstruct 4D content from the image matrix produced in the first stage.
@@ 3.1 Image matrix generation
In this stage, our goal is to generate dense multi-frame multi-view images for reconstruction, which can be denoted as a matrix of image
@@
where $V$ is the number of views, $F$ is the number of video frames, and $(H,W)$ is the size of images. We aim to construct a generative model that allows us to directly sample $\mathcal{I}\sim p(\mathcal{I})$.
Now, let us first distract our attention to reviewing existing diffusion-based generators for video and multi-view images, which can be utilized for sampling realistic images through the following probabilistic flow ODE:
@@
Here, $\mathrm{x}=\left\{I_{i}\in\mathbb{R}^{H\times W\times 3}\right\}_{i=1}^{N}$ is a series of images with $N$ frames or $N$ views, $\nabla_{\mathrm{x}}\log p\left(\mathrm{x};\sigma\right)$ is the score function, which can be parameterized as $\nabla_{\mathrm{x}}\log p\left(\mathrm{x};\sigma\right)\approx\left(D_{\theta}(\mathrm{x};\sigma)\right)/\sigma^{2}$ karras2022elucidating ; blattmann2023stable , where $D_{\theta}(\mathrm{x};\sigma)$ is a neural network trained via denoising score matching.
We want to extend the above formulation to the sampling of $\mathcal{I}$. The question is, how do we estimate the score function of the joint distribution of $V\times F$ images?
For simplicity, let $\mathcal{I}_{-i,j}\triangleq\{I_{i^{\prime},j}|1\leq i^{\prime}\leq V,i^{\prime}\neq i\}$, $\mathcal{I}_{i,-j}\triangleq\{I_{i,j^{\prime}}|1\leq j^{\prime}\leq F,j^{\prime}\neq j\}$ and $\mathcal{I}_{-i,-j}\triangleq\{I_{i^{\prime},j^{\prime}}|1\leq i^{\prime}\leq V,1\leq j^{\prime}\leq F,i^{\prime}\neq i,j^{\prime}\neq j\}$. We first make an assumption on the structure of $p(\mathcal{I})$.
@@
Given any image $I_{i,j}$, the underlying geometry $\mathcal{I}_{-i,j}$ and the dynamics $\mathcal{I}_{i,-j}$ are conditionally independent, i.e.,
$p\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{I}_{-i,j}},{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\mathcal{I}_{i,-j}}|I_{i,j}\right)=p\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{I}_{-i,j}}|I_{i,j}\right)p\left({\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\mathcal{I}_{i,-j}}|I_{i,j}\right).$ (3)
The assumption 3.1 implies that given the front view of a 3D object, its motion as seen from the front does not correlate with its appearance from the back, which aligns with our intuition. A natural corollary is that the mollified distribution derived by adding Gaussian noise into the data distribution still maintains conditional independence:
@@
where $\alpha\in\mathbb{R}$ is a constant and $\varepsilon_{i,j}\in\mathbb{R}^{H\times W\times 3}$ are independent Gaussian noises. Then we have
$p\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\mathcal{I}}_{-i,j}},{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\hat{\mathcal{I}}_{i,-j}}|\hat{I}_{i,j}\right)=p\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\mathcal{I}}_{-i,j}}|\hat{I}_{i,j}\right)p\left({\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\hat{\mathcal{I}}_{i,-j}}|\hat{I}_{i,j}\right).$ (5)
This nice property allows us to sample the desired image matrix by progressively denoising from pure Gaussian noise through the combination of two estimated scores of its marginal distribution, which can be obtained from the pretrained video and multi-view diffusion models respectively. Therefore, we can derive our main theorem.
@@
$p\left(\hat{\mathcal{I}}\right)$ $=p\left(\hat{I}_{i,j},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\mathcal{I}}_{-i,j}},{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\hat{\mathcal{I}}_{i,-j}},{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\hat{\mathcal{I}}_{-i,-j}}\right)$ (7)
$=p\left(\hat{I}_{i,j},{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\hat{\mathcal{I}}_{-i,-j}}|{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\mathcal{I}}_{-i,j}},{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\hat{\mathcal{I}}_{i,-j}}\right)p\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\mathcal{I}}_{-i,j}},{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\hat{\mathcal{I}}_{i,-j}}\right).$ (8)
Note that for any $\hat{I}_{i^{\prime},j^{\prime}}\in{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\hat{\mathcal{I}}_{-i,-j}}$, $I_{i^{\prime},j^{\prime}}$ and $I_{i,j}$ are independent conditioned on $I_{i^{\prime},j}$ by corollary 3.1, hence

### v3

@@ Abstract
Recent advancements in 3D generation are predominantly propelled by improvements in 3D-aware image diffusion models which are pretrained on Internet-scale image data and fine-tuned on massive 3D data, offering the capability of producing highly consistent multi-view images. However, due to the scarcity of synchronized multi-view video data, it is impractical to adapt this paradigm to 4D generation directly. Despite that, the available video and 3D data are adequate for training video and multi-view diffusion models separately that can provide satisfactory dynamic and geometric priors respectively. To take advantage of both, this paper present Diffusion2, a novel framework for dynamic 3D content creation that reconciles the knowledge about geometric consistency and temporal smoothness from these models to directly sample dense multi-view multi-frame images which can be employed to optimize continuous 4D representation. Specifically, we design a simple yet effective denoising strategy via score composition of pretrained video and multi-view diffusion models based on the probability structure of the target image array. Owing to the high parallelism of the proposed image generation process and the efficiency of the modern 4D reconstruction pipeline, our framework can generate 4D content within few minutes. Additionally, our method circumvents the reliance on 4D data, thereby having the potential to benefit from the scaling of the foundation video and multi-view diffusion models. Extensive experiments demonstrate the efficacy of our proposed framework and its ability to flexibly handle various types of prompts.
@@ 1 Introduction
Figure 1: Diffusion2 is designed to generate dynamic content by creating a dense multi-frame multi-view image matrix in a highly parallel denoising diffusion process with the combination of the foundation video diffusion model and multi-view diffusion model. The generated image can be used to construct a full 4D representation by feeding them into an off-the-shelf 4D reconstruction pipeline.
@@
Spurred by the advances from generative image models ho2020denoising ; song2020denoising ; song2020score ; karras2022elucidating ; zhang2023adding , automatic 3D content creation poole2022dreamfusion ; wang2024prolificdreamer ; tang2023dreamgaussian ; hong2023lrm has witnessed remarkable progress in terms of efficiency, fidelity, diversity, and controllability. Coupled with the breakthroughs in 4D representation yang2023deformable ; wu20234d ; duan20244d , these advances further foster substantial development in dynamic 3D content generation singer2023text ; bahmani20234d ; jiang2023consistent4d ; zhao2023animate124 ; ren2023dreamgaussian4d ; gao2024gaussianflow , which holds significant value across a wide range of applications in animation, film, game, and MetaVerse.
@@
Recently, 3D content generation has achieved considerable breakthroughs in efficiency. Some works liu2023zero ; liu2023syncdreamer ; shi2023mvdream ; wang2023imagedream ; tang2024mvdiffusion++ inject stereo knowledge into the image generation model, enabling these 3D-aware image generators to produce consistent multi-view images, thereby effectively stabilizing and accelerating the optimization. Other efforts hong2023lrm ; chen2024v3d ; voleti2024sv3d ; zuo2024videomv ; tang2024lgm attempt to directly generate 3D representations, such as triplane chen2022tensorf or Gaussian splatting kerbl3Dgaussians . However, the efficiency improvement from these works is largely data-driven wu2023omniobject3d ; reizenstein2021common ; yu2023mvimgnet ; deitke2023objaverse . Consequently, it is infeasible to adapt them to 4D generation due to the scarcity of synchronized multi-view video data. Therefore, most existing 4D generation works jiang2023consistent4d ; yin20234dgen ; ren2023dreamgaussian4d still adopt the score distillation sampling (SDS) approach and suffer from slow or unstable optimization.
@@
However, despite the paucity of 4D data, there are vast available monocular video data and static multi-view image data. Existing works have demonstrated that it is feasible to train diffusion-based generative models learning the distribution of these two classes of data independently  voleti2024sv3d ; liu2023syncdreamer ; tang2024mvdiffusion++ ; blattmann2023align ; blattmann2023stable . Considering that video diffusion model stores the prior of motion and temporal smoothness, and multi-view diffusion model has sound knowledge of geometrical consistency, combining the two generative models to generate 4D assets becomes a highly promising and appealing approach.
@@
To leverage both of them, inspired by the previous work bar2023multidiffusion ; lee2023syncdiffusion , we propose a novel 4D generation framework that reconciles the video and multi-view diffusion priors to directly sample multi-frame multi-view image arrays imitating the photographing process of 4D object. To demonstrate how we realize this combination, we assume that such an image matrix $\mathcal{I}$ has a nice structure: the elements $\mathcal{I}_{i,j}$ and $\mathcal{I}_{k,l}$ not in the same row and column ($i\neq k,j\neq l$) are conditionally independent of each other given $\mathcal{I}_{i,l}$ or $\mathcal{I}_{k,j}$. Based on this property, we design a simple yet effective denoising sampling strategy in which the estimated score is just the convex combination of the scores predicted by two foundation diffusion models. Our formulation is easy to adapt to various prompts including reference image, single-view video, and static 3D content. Unlike the existing SDS-based counterparts, Diffusion2 is highly parallel. Combined with the efficient 4D reconstruction pipeline, our method can generate high-fidelity and diverse 4D assets within just a few minutes. Besides, our approach can also potentially benefit from the further advancement of foundation diffusion models videoworldsimulators2024 .
@@
Our contributions are threefold: (i) We present a novel 4D content generation approach that achieves zero-shot generation of realistic multi-view multi-frame image arrays, which can be integrated with an off-the-shelf modern 4D reconstruction pipeline to efficiently create 4D content. (ii) We identify the conditional independence within elements constituting the image arrays. Based on it, we devise a simple yet effective joint denoising strategy that combines video diffusion model and multi-view diffusion model to directly sample multi-view multi-frame images from their natural joint distribution. (iii) Systematic experiments demonstrate that our proposed method can achieve satisfactory results under different types of prompts, including single image, single-view video, and static 3D content.
2 Related work
@@
3D generation   3D generation aims at creating static 3D content from different type of conditions like text or reference image. Early efforts employed GAN-based approaches gao2022get3d ; schwarz2020graf . Recently, significant breakthroughs have been achieved with diffusion models ho2020denoising . DreamFusion poole2022dreamfusion introduced score distillation sampling (SDS) to unleash the creativity in diffusion models. A series of subsequent works wang2024prolificdreamer ; shi2023mvdream ; wang2023imagedream ; pan2024enhancing ; tang2023dreamgaussian ; yi2023gaussiandreamer continuously address challenges such as multi-face Janus issues and slow optimization. On the other hand, with the development of large scale 3D datasets yu2023mvimgnet ; deitke2023objaverse , many works try to directly generate 3D contents by using diffusion models. Some studies nichol2022point ; jun2023shap ; hong2023lrm ; tang2024lgm ; wang2024crm have explored the direct generation of 3D representations. Another line of research liu2023zero ; liu2023syncdreamer ; long2023wonder3d ; chen2024v3d ; tang2024mvdiffusion++ focuses on generating multi-view images with sufficient 3D consistency to be used for reconstruction. We also adopt this approach of directly generating consistent images for reconstruction. But unlike the above 3D counterparts, there is no large-scale multi-view synchronized video data. Therefore, we opt to combine geometrical consistency priors and video dynamic priors to generate images.
@@
Video generation   Video generation is an active field that has gained increasing popularity. Recent diffusion-based methods have exhibited unprecedented levels of realism, diversity, and controllability. Video LDM blattmann2023align is the pioneer work to apply the latent diffusion framework rombach2022high to video generation. The subsequent work SVD blattmann2023stable followed its architecture and made effective improvements to the training recipe. W.A.L.T gupta2023photorealistic employed a transformer with window attention tailored for spatiotemporal generative modeling to generate high-resolution videos. VDT lu2023vdt introduced the video diffusion transformer and a spatial-temporal mask mechanism to flexibly capture long-distance spatiotemporal context. The recently introduced SORA videoworldsimulators2024 demonstrated a remarkable capability to generate long videos with intuitively physical fidelity. Models trained on large-scale video data can generate videos with realistic dynamics. Besides, video diffusion models can also provide effective prior for 3D generators chen2024v3d ; han2024vfusion3d ; voleti2024sv3d . Therefore, we build our method on this flourishing domain.
@@
4D generation   Animating category-agnostic objects is challenging and has drawn considerable attention from both academia and industry. Unlike 3D generation, 4D generation requires both consistent geometry and realistic dynamics. Recent works on this domain can be categorized based on the input condition. Some of them create 4D content from text or single image. For instance, MAV3D singer2023text first employs SDS in text-to-4D tasks. 4D-fy bahmani20234d hybridizes different diffusion priors during SDS training. But they suffer from extremely slow generation. DreamGaussian4D ren2023dreamgaussian4d adopts deformable 3D Gaussian yang2023deformable as the underlying 4D representation and exports mesh for texture refinement, significantly improving efficiency. Another line of work predicts dynamic objects from a single-view video with largely dictated motion. Consistent4D jiang2023consistent4d proposes to use SDS approach for geometry consistency and frame interpolation loss for temporal continuity. 4DGen yin20234dgen further grounds the 4D content creation with pseudo labels. Efficient4D pan2024fast mimics a photogrammetry-based neural volumetric video reconstruction pipeline by directly generating multi-view videos for reconstruction. Although we share the same SDS-free design, Efficient4D possesses a strong architectural preference for the fundamental diffusion model and is incapable of synthesizing novel dynamics. Compared to previous works, our framework can efficiently generate diverse 4D contents from different input prompts, avoiding the slow and unstable optimization. In addition, our method has the potential to continuously benefit from the scalability of underlying diffusion models.
@@ 3 Method
As depicted in Figure 2, the proposed 4D generation framework Diffusion2 adopts a two-stage pipeline: dense observation synthesis followed by reconstruction. In Section 3.1 (stage-1), we will elaborate on how to generate dense multi-view multi-frame images for reconstruction through a highly parallelizable denoising process by reconciling the pretrained video diffusion prior and multi-view diffusion prior, and why it is feasible. In Section 3.2 (stage-2), we will briefly introduce how to robustly reconstruct 4D content from the images produced in the first stage.
@@ 3.1 Image matrix generation
Figure 2: The overall pipeline of Diffusion2. (i) Given a reference image, Diffusion2 first independently generates the animation under the reference view (denoted $\mathcal{I}_{1:V,0}$) and the multi-view images at the reference time (denoted $\mathcal{I}_{0,1:F}$) as the condition for the subsequent generation of the full matrix, denoted $\mathcal{I}$. Depending on the type of given prompt, the condition images $\mathcal{I}_{1:V,0}$ or $\mathcal{I}_{0,1:F}$ can be specified by users. (ii) Then, Diffusion2 directly samples a dense multi-frame multi-view image array by blending the estimated scores with a weighting scheduler from pretrained video and multi-view diffusion models in the reverse diffusion process. (iii) The generated image arrays are employed as supervision to optimize a continuous 4D content representation.
In this stage, our goal is to generate dense multi-frame multi-view images for reconstruction, which can be denoted as a matrix of images:
@@
where $V$ is the number of views, $F$ is the number of video frames, and $(H,W)$ is the size of images. We aim to construct a generative model that allows us to directly sample natural $\mathcal{I}\sim p(\mathcal{I})$.
Now, let us first divert our focus to reviewing existing diffusion-based generators for video and multi-view images, which can be utilized for sampling realistic image sequence through the probabilistic flow ODE as below:
@@
Here, $\mathrm{x}=\left\{I_{i}\in\mathbb{R}^{H\times W\times 3}\right\}_{i=1}^{N}$ is a series of images with $N$ frames or $N$ views, $\nabla_{\mathrm{x}}\log p\left(\mathrm{x};\sigma\right)$ is the score function, which can be estimated as $\nabla_{\mathrm{x}}\log p\left(\mathrm{x};\sigma\right)\approx\left(D_{\theta}(\mathrm{x};\sigma)\right)/\sigma^{2}$ karras2022elucidating ; blattmann2023stable , where $D_{\theta}(\mathrm{x};\sigma)$ is a neural network trained via denoising score matching.
We want to extend the above formulation to the sampling of $\mathcal{I}$. The question is, how do we estimate the score of the joint distribution of $V\times F$ images?
For clarity, denote $\mathcal{I}_{-i,j}\triangleq\{I_{i^{\prime},j}|1\leq i^{\prime}\leq V,i^{\prime}\neq i\}$, $\mathcal{I}_{i,-j}\triangleq\{I_{i,j^{\prime}}|1\leq j^{\prime}\leq F,j^{\prime}\neq j\}$ and $\mathcal{I}_{-i,-j}\triangleq\{I_{i^{\prime},j^{\prime}}|1\leq i^{\prime}\leq V,1\leq j^{\prime}\leq F,i^{\prime}\neq i,j^{\prime}\neq j\}$. We first make an assumption on the structure of $p(\mathcal{I})$.
@@
Given any image $I_{i,j}$, the underlying geometry $I_{i^{\prime},j}\in\mathcal{I}_{-i,j}$ and the dynamics $I_{i,j^{\prime}}\in\mathcal{I}_{i,-j}$ are conditionally independent, i.e.,
$p\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}I_{i^{\prime},j}},{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}I_{i,j^{\prime}}}|I_{i,j}\right)=p\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}I_{i^{\prime},j}}|I_{i,j}\right)p\left({\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}I_{i,j^{\prime}}}|I_{i,j}\right).$ (3)
The assumption 3.1 implies that, given the front view at now of a 3D object, what its front view looks like in the future (or past) does not correlate with what the back view looks like now, which aligns with our intuition. Actually, assumption 3.1 is too strong to hold for some specific cases, yet this does not violate the correctness of our main theorem. For the conciseness and readability of the main text, we leave the discussion of the scope of the assumption and a more rigorous proof in Section A.3.
A natural corollary is that the mollified distribution derived by adding Gaussian noise into the data distribution still maintains conditional independence:
@@
where $\alpha\in\mathbb{R}$ is a constant and $\varepsilon_{i,j}\in\mathbb{R}^{H\times W\times 3}$ are independent Gaussian noises. Then equation (3) also holds for $\hat{\mathcal{I}}$.
This nice property allows us to sample the desired image matrix by progressively denoising from pure Gaussian noise using the combination of two estimated scores for its column and row, which can be obtained from the pretrained video and multi-view diffusion models respectively. Therefore, we can establish our main theorem.
@@
$p\left(\hat{\mathcal{I}}\right)$ $=p\left(\hat{I}_{i,j},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\mathcal{I}}_{-i,j}},{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\hat{\mathcal{I}}_{i,-j}},{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\hat{\mathcal{I}}_{-i,-j}}\right)=p\left(\hat{I}_{i,j},{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\hat{\mathcal{I}}_{-i,-j}}|{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\mathcal{I}}_{-i,j}},{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\hat{\mathcal{I}}_{i,-j}}\right)p\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\mathcal{I}}_{-i,j}},{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\hat{\mathcal{I}}_{i,-j}}\right).$ (6)
Note that $\forall\hat{I}_{i^{\prime},j^{\prime}}\in{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\hat{\mathcal{I}}_{-i,-j}}$, $I_{i^{\prime},j^{\prime}}$ and $I_{i,j}$ are independent conditioned on $I_{i^{\prime},j}$ by corollary 3.1, hence